# Agentic AI Security: Threats, Defenses, Evaluation, and Open Challenges

### v1

@@ 1 Introduction
Agentic AI represents the next stage in this natural evolution of AI systems powered by LLMs and other Generative AI models. Agentic AI systems are characterized by autonomy, goal-directed reasoning, planning, and the ability to act upon digital or physical environments through tools, APIs, or robotic embodiments [15, 16]. Unlike static LLMs, agentic systems maintain persistent memory, deliberate across time, coordinate with other agents, and adapt dynamically to changing contexts. These capabilities position agentic AI as a powerful general-purpose automation platform rather than a reactive model that depends on continuous human input for task completion.
@@ 2 Motivation and Contributions
A number of these security problems in agentic AI systems compound upon the vulnerabilities and deficiencies of the base LLMs [36], while others are novel and occur due to the unique landscape of agent-agent interactions. Prompt injection attacks remains a critical LLM vulnerability, permitting adversaries to manipulate agent behavior through crafted inputs [37, 38, 39]. Lupinacci et al. demonstrated that 94.4% of state-of-the-art LLM agents are vulnerable to prompt injection, 83.3% to retrieval-based backdoors, and 100% to inter-agent trust exploits [40]. Further, researchers at Anthropic observed that generative models, when given directive autonomy, engaged in misaligned behaviors such as blackmail or corporate espionage to fulfill goals, even when those behaviors diverged from human ethical standards [41]. It is evident that these security vulnerabilities can result in highly adverse consequences for some of the proposed critical application areas for agentic AI frameworks, such as healthcare.
@@
Given the rapid acceleration of progress in developing and utilizing agentic AI systems, it is paramount for the community to understand and address the aforementioned security risks. However, owing to the large body of work in the area, it is important to systemize current research efforts and create a taxonomy of all the security threats posed by agentic AI. Furthermore, for truly securing these agentic systems, it is equally consequential to understand the state-of-the-art in defending against known attacks/threats, and provide an overview of open challenges that remain. Through this survey, we seek to make significant progress along each of these directions, and our aim is to galvanize research efforts in developing safe and secure agentic AI frameworks, thereby accelerating their adoption in society. Our work is also distinct from existing surveys in this space, and tackles an overlooked problem. Existing surveys on autonomous AI agents [49, 50, 51, 52, 53] mostly detail their capabilities and propose evaluation benchmarks, while others explore topics such as trust and risk management [54] without a singular focus on security, unlike our work. On the other hand, risk management frameworks such as the NIST Generative AI Profile [55] provide baseline guidance, but adapting them to autonomous agents remains a work in progress.
@@ 3.1.1 Direct and Indirect Prompt Injection
Figure 3: An example showcasing unintentional prompt injection.
@@ 3.3.2 Threat Actor Perspective
From the perspective of a threat actor, multi-agent systems spanning multiple organizational domains provide significant opportunities to conceal responsibility for malicious actions. Each domain typically enforces independent logging, retention, and auditing policies, preventing a unified trace of activity [109]. Once input data is processed by an LLM, it is transformed into distributed latent representations, eliminating persistent identifiers that could link actions to their source [142]. Unlike conventional software systems that can implement taint checking or explicit information flow tracking [143], these latent representations make tracing infeasible [144]. Adversaries can exploit this by compromising an agent in Domain A, injecting deceptive but plausible instructions, and indirectly triggering harmful actions in Domain B, all while masking their identity and intent [145, 146, 147]. Even with auditing at the agent level, inter-agent relationships and cross-domain causality remain hidden, significantly complicating accountability. Current interpretability tools, such as influence functions [chhabra2024data, chhabraoutlier, 148, 149], activation tracing [150], or model-source attribution queries [151, 152, 153], provide limited visibility and often fail in these complex multi-agent settings.
@@ 4 Defenses and Security Controls
To safeguard agentic AI systems from threats, various defense strategies and frameworks have been proposed. However, due to the constant evolution of attack vectors and threats, defense methods need to be continuously optimized and evaluated. To facilitate progress in developing better agentic AI defense mechanisms, we now discuss existing and current approaches. We also delineate potential shortcomings of these approaches.
@@ 4.1 Prompt-Injection-Resistant Designs
Figure 6: Some Defenses Against Prompt Injection Attacks.
@@ 4.1.1 Agent-Focused
The authors in [176] propose an instruction hierarchy, which establishes priority levels for different instruction sources so that user-provided instructions are always prioritized over potentially malicious instructions embedded in retrieved content. Other works introduce adversarial training methods that teach the model to resist prompt injection attacks [59]. Complementary to these efforts, research has also been undertaken to utilize circuit breaking or task drifting approaches to recognize and reject adversarial patterns while preserving intended functionality [178, 150].
@@ 4.1.3 System-Focused
System-level defenses aim to safeguard LLM agents by integrating external verification, control mechanisms, and constrained execution environments [59]. However, designing agent systems inherently robust to attacks remains a challenging technical problem. Analogous to defending against traditional ML/AI adversarial examples in computer vision, completely preventing prompt injection is an open problem [186]. Despite this issue, recent works have introduced design patterns for system-focused protection, such as Action-Selector, Plan-then-Execute, LLM Map-Reduce, Code-Then-Execute, Dual-LLM, and Context-Minimization patterns [59]. Other system-focused defenses, such as Melon, employ constrained execution sandboxes and verification loops to limit the impact of potentially malicious instructions on downstream systems, serving as a defense against IPI attacks [72]. We will now discuss some sub-categories of system-focused defenses.
@@ 4.1.4 Training-based and Training-free Defenses
Figure 7: Policy Filtering and Enforcement Defense Strategies.
@@ 5 Benchmarks and Evaluation
To assess the impact of various agentic AI security vulnerabilities associated we discussed previously (as well the potency of defense strategies) it is necessary to undertake system evaluation via robust benchmarks. The first benchmarks proposed for agentic AI primarily focused on competence and sought to study whether an agent could complete a specified task under controlled conditions. Lately, as deployments edge closer to production, evaluation has shifted toward reliability, safety, and control. Thus, in this section, we cover: (i) landscape benchmarks used to gauge general capability, (ii) security–specific frameworks that probe failure modes such as harmful misuse, and (iii) methodological advances that improve fidelity, comparability, and reproducibility. Table 1 summarizes existing benchmarks spanning both capability and security-specific focuses, highlighting their domains, safety focus, and evaluation methodologies.
5.1 Landscape benchmarks for capability
[…]
5.2 Security-Specific and safety-specific benchmarks
@@
ST-WebAgent-Bench [231] Enterprise CuP + Risk Policy compliance
@@
OS-Harm [232] Computer LLM Safety + accuracy Desktop effects
@@
Annotations: $\checkmark$ = fully supported; - = not applicable/not supported; $\sim$ = varies by implementation or framework-dependent.
Columns: Safety Focus = explicit evaluation of security/safety risks; Process-Aware = trajectory-level evaluation beyond end-state metrics; Multi-Turn = support for extended agent interactions; Sandboxed = isolated/controlled environment for safe testing; Judge Type = evaluation method (Rule-based: deterministic scoring, LLM: model-based evaluation, blank: not specified); Reliability Metric = method for measuring consistency/robustness.

### v2

@@ 1 Introduction
Agentic AI represents the next stage in this natural evolution of AI systems powered by LLMs and other Generative AI models. Agentic AI systems are characterized by autonomy, goal-directed reasoning, planning, and the ability to act upon digital or physical environments through tools, APIs, or robotic embodiments [120, 163]. Figure 1 illustrates the general interaction flows of such systems. Unlike static LLMs, agentic systems maintain persistent memory, deliberate across time, coordinate with other agents, and adapt dynamically to changing contexts. These capabilities position agentic AI as a powerful general-purpose automation platform rather than a reactive model that depends on continuous human input for task completion.
Figure 1: Agentic AI architecture: general components and interaction flows. The agent performs planning and reasoning, retrieves and stores information in memory, executes actions via tools, and interacts with both the user and the external environments to accomplish multi-step tasks autonomously.
@@ 2 Motivation and Contributions
A number of these security problems in agentic AI systems compound upon the vulnerabilities and deficiencies of the base LLMs, while others are novel and occur due to the unique landscape of agent-agent interactions. Prompt injection attacks remains a critical LLM vulnerability, permitting adversaries to manipulate agent behavior through crafted inputs [258, 123]. Lupinacci et al. demonstrated that 94.4% of state-of-the-art LLM agents are vulnerable to prompt injection, 83.3% to retrieval-based backdoors, and 100% to inter-agent trust exploits [142]. Further, researchers at Anthropic observed that generative models, when given directive autonomy, engaged in misaligned behaviors such as blackmail or corporate espionage to fulfill goals, even when those behaviors diverged from human ethical standards [10]. It is evident that these security vulnerabilities can result in highly adverse consequences for some of the proposed critical application areas for agentic AI frameworks, such as healthcare.
@@
Given the rapid acceleration of progress in developing and utilizing agentic AI systems, it is paramount for the community to understand and address the aforementioned security risks. However, owing to the large body of work in the area, it is important to systemize current research efforts and create a taxonomy of all the security threats posed by agentic AI. Furthermore, for truly securing these agentic systems, it is equally consequential to understand the state-of-the-art in defending against known attacks/threats, and provide an overview of open challenges that remain. Through this survey, we seek to make significant progress along each of these directions, and our aim is to galvanize research efforts in developing safe and secure agentic AI frameworks, thereby accelerating their adoption in society.
Remark. Our work is also distinct from existing surveys in this space, and tackles an overlooked problem. Existing surveys on autonomous AI agents [192, 151, 132, 77, 3] mostly detail their capabilities and propose evaluation benchmarks, while others explore topics such as trust and risk management [180] without a singular focus on security, unlike our work. On the other hand, risk management frameworks such as the NIST Generative AI Profile [5] provide baseline guidance, but adapting them to autonomous agents remains a work in progress.
@@ 3 Taxonomy of Security Threats
Figure 2: Taxonomy of Agentic AI Security Threats.
Attacks and threats can be multifaceted, and their behavior often spans multiple threat patterns. Our taxonomy is therefore not meant to define mutually exclusive categories; rather, it organizes attacks according to their primary distinguishing properties for analytical clarity.
@@ 3.1.2 Intentional and Non-Intentional Prompt Injection
Figure 4: An example showcasing unintentional prompt injection.
@@ 3.3.1 Protocol-level MCP and A2A attacks
As illustrated in Figure 6(a), an agent may issue a request to retrieve the contact information of a business from the internet using MCP protocol. An attacker can exploit a leaked or compromised response token to tamper with the returned contact details. Because the agent treats the token as valid, it subsequently trusts and acts upon the manipulated information, leading to insecure downstream behavior. On the other hand, if an attacker gains access to the database or file system consulted by the agent through MCP protocol, they can directly poison the underlying data, thereby influencing the agent’s behavior when it consumes and acts upon this compromised information.
@@
In a multi-agent system where the responsibilities are divided, as depicted in Figure 6(b), when an existing agent issues a discovery request to identify an agent for email summarization, an adversary can inject a malicious agent by forging a discovery response through the A2A protocol. Similarly, an attacker may hijack A2A-based interactions to induce other agents to store malicious websites as trusted user preferences (e.g., a favorite shopping site), thereby influencing subsequent agent decisions and actions.
@@ 3.3.2 Threat Actor Perspective
From the perspective of a threat actor, multi-agent systems spanning multiple organizational domains provide significant opportunities to conceal responsibility for malicious actions. Each domain typically enforces independent logging, retention, and auditing policies, preventing a unified trace of activity [116]. Once input data is processed by an LLM, it is transformed into distributed latent representations, eliminating persistent identifiers that could link actions to their source [86]. Unlike conventional software systems that can implement taint checking or explicit information flow tracking [68], these latent representations make tracing infeasible [203]. Adversaries can exploit this by compromising an agent in Domain A, injecting deceptive but plausible instructions, and indirectly triggering harmful actions in Domain B, all while masking their identity and intent [168, 204, 58]. Even with auditing at the agent level, inter-agent relationships and cross-domain causality remain hidden, significantly complicating accountability. Current interpretability tools, such as influence functions [47, 46], activation tracing [1], or model-source attribution queries [49, 170, 190], provide limited visibility and often fail in these complex multi-agent settings.
@@ 4 Defenses and Security Controls
To safeguard agentic AI systems from threats, various defense strategies and frameworks have been proposed. However, due to the constant evolution of attack vectors and threats, defense methods need to be continuously optimized and evaluated. To facilitate progress in developing better agentic AI defense mechanisms, we now discuss existing and current approaches. We also delineate potential shortcomings of these approaches. Moreover, we present Table 1, which shows secure-by-design coverage of security-critical defense components across defense approaches.
@@ 4.1 Prompt-Injection-Resistant Designs
Figure 7: Defenses Against Prompt Injection Attacks: (a) Agent-Focused, (b) User-Focused, and (c) System-Focused.
@@ 4.1.1 Agent-Focused
The authors in [223] propose an instruction hierarchy, which establishes priority levels for different instruction sources so that user-provided instructions are always prioritized over potentially malicious instructions embedded in retrieved content. Other works introduce limitations of executable actions by agent to resist prompt injection attacks such as Action-Selector and Plan-then-Execute design patterns [17]. Complementary to these efforts, research has also been undertaken to utilize circuit breaking or task drifting approaches to recognize and reject adversarial patterns while preserving intended functionality [273, 1].
@@ 4.1.3 System-Focused
System-level defenses aim to safeguard LLM agents by integrating external verification, control mechanisms, and constrained execution environments [17]. However, designing agent systems inherently robust to attacks remains a challenging technical problem. Analogous to defending against traditional ML/AI adversarial examples in computer vision, completely preventing prompt injection is an open problem [209]. Despite this issue, recent works have introduced design patterns for system-focused protection, such as LLM Map-Reduce, Code-Then-Execute, Dual-LLM, and Context-Minimization patterns [17]. With a similar view, TrustAgent [91] introduces an Agent-Constitution–guided framework that enforces safety in LLM-based agent planning through pre-, in-, and post-planning safeguards, demonstrating improved safety and helpfulness across domains. Other system-focused defenses, such as Melon, employ training‑free runtime detection approach that performs masked re‑execution and tool‑call similarity checks and verification loops to limit the impact of potentially malicious instructions on downstream systems, serving as a defense against IPI attacks [268]. We will now discuss some sub-categories of system-focused defenses.
@@
Trade-off comparison of agent-, user-, and system‑focused defenses. In summary, each defense category presents distinct trade-offs between robustness, automation, and deployment cost. User-focused defenses provide strong guarantees through explicit human verification, yet sacrifice scalability and automation, making them unsuitable for autonomous agents [238, 17]. They also fail when the attacker is the user themselves. In contrast, offering seamless protection without human intervention, agent-focused defenses internalize security within the model but often require retraining and may degrade general-purpose capabilities or fail against adaptive attacks [105]. System-focused defenses strike a middle ground by preserving model generality while enforcing external controls and verification; however, they increase system complexity and remain vulnerable to sophisticated multi-stage or multi-agent based attacks [197, 209]. Consequently, practical deployments increasingly favor hybrid designs that combine complementary defenses across these categories rather than relying on a single strategy.
@@ 4.1.4 Training-based and Training-free Defenses
Figure 8: Policy Filtering and Enforcement Defense Strategies: (a) Governance-Centric, (b) Signal-Centric.
@@ 4.5 Standards and Organizational Measures
Table 1: Secure-by-Design Defense Mechanisms Coverage Across Agentic AI Security Approaches.
Defense Approach Input Governance Tool-Call Guards Policy Engine Sandboxing/ Isolation Rate Limiting Audit Trails Rollback/ Recovery
Prompt-Injection-Resistant Designs
PromptArmor [197]
StruQ [34]
SecAlign [36]
Sign-Prompt [206]
Attn. Tracker [97]
JATMO [171]
TaskShield [103]
Melon [268]
Policy Filtering & Enforcement
LlamaGuard [99]
LlavaGuard [88]
GuardAgent [239]
AgentSpec [226]
ShieldAgent [41]
R²-Guard [110]
Sandboxing & Capability Confinement
SandboxEval [174]
IsolateGPT [238]
ToolEmu [188]
Monitoring & Adaptive Systems
Pro2Guard [225]
Arms Race [217]
Dec. RE [176]
CB [273]
Input Governance: Input validation, sanitization, and prompt injection detection mechanisms to prevent malicious inputs from reaching the agent. Tool-Call Guards: Runtime verification and constraints on tool/API invocations to prevent unauthorized or dangerous actions. Policy Engine: Formal policy specification and enforcement mechanisms to ensure agent behavior aligns with safety requirements. Sandboxing/Isolation: Isolated execution environments that contain potential damage from risky operations. Rate Limiting: Request throttling and resource usage controls to prevent denial-of-service and resource exhaustion. Audit Trails: Logging and traceability of agent actions for accountability and post-incident analysis. Rollback/Recovery: Mechanisms to reverse harmful actions or restore the system to safe states after detection of violations.
Annotations = Explicitly supported; = Not addressed or not applicable.
@@ 5 Benchmarks and Evaluation
To assess the impact of various agentic AI security vulnerabilities associated we discussed previously (as well the potency of defense strategies) it is necessary to undertake system evaluation via robust benchmarks. The first benchmarks proposed for agentic AI primarily focused on competence and sought to study whether an agent could complete a specified task under controlled conditions. Lately, as deployments edge closer to production, evaluation has shifted toward reliability, safety, and control. Consequently, recently DoomArena [19] has introduced security-oriented evaluation frameworks, a modular and configurable platform that enables cross-environment attack composition and fine-grained threat modeling for systematically probing agent failure modes and defense limitations. Thus, in this section, we cover: (i) landscape benchmarks used to gauge general capability, (ii) security–specific frameworks that probe failure modes such as harmful misuse, and (iii) methodological advances that improve fidelity, comparability, and reproducibility. Table 2 summarizes existing benchmarks spanning both capability and security-specific focuses, highlighting their domains, safety focus, and evaluation methodologies.
@@
VisualWebArena [117] Web $\sim$ End-state, fuzzy Visually grounded
AndroidWorld [179] Android Rule-based End-state Dynamic parameterization
WebShop [247] E-commerce Rule-based End-state, Programmatic Reward Scalable
AgentBench [134] Multi-domain Rule-based End-state, step-wise Eight environments
WCA [150] Web $\sim$ End-state, fuzzy Labor-intensive "chores"
CRAB [242] Multi-domain Rule-based Graph-based Graph-based evaluation
AssistantBench [252] Open web $\sim$ Realistic & time-consuming
@@
ST-WebAgent-Bench [125] Enterprise CuP, Risk Policy compliance
@@
OS-Harm [8] Computer LLM Safety, Accuracy Desktop effects
@@
AgentDojo [61] Tool-use Rule-based Utillity, ASR Dynamic Framework
InjecAgent [258] Tool-use Rule-based ASR Indirect Prompt Injection
ASB [260] Multi-domain Rule-based ASR, RR, BP, PNA, NRP Multi-scenario
PrivacyLens [195] Personalized Comm $\sim$ LR, Helpfulness Privacy norm awareness
Agent-SafetyBench [264] Tool-use LLM Safety Score, Helpfulness Comprehensive
SAB [251] Embodied AI $\sim$ Rejection Rate, Risk Rate Embodied agents in simulated environment
CAIBench [191] Cybersecurity Rule-based $\sim$ Simultaneous offense and defense evaluation
WASP [69] Web $\sim$ ASR (intermediate & end-to-end), Utility Security by incompetence
SafeArena [219] Web $\sim$ TSR, RR, NSS, ARIA Deliberate misuse focused
OAS [222] Multi-domain $\sim$ FR, DR Multi-user dynamics
MCP-SafetyB [272] MCP Rule-based TSR, ASR Real MCP servers
MCP-SecB [246] MCP Rule-based ASR, RR, PSR Real MCP servers
MAGPIE [108] Multi-domain $\sim$ LR, TSR (Constraints), MR, Consensus Multi-agent
[…]
Full forms: Leakage Rate (LR), Misclassification Rate (MR), Refusal Rate (RR), Protection Success Rate (PSR), Task Success Rate (TSR), Failure Rate(FR), Disagreement Rate (DR), Normalized Safety Score (NSS), ARIA risk levels, Performance under No Attack (PNA), Benign performance (BP), and Net Resilient Performance (NRP).
5.1 Landscape benchmarks for capability
[…]
5.2 Security-Specific and safety-specific benchmarks
@@
Red-team corpora, dynamic web testbeds, and OS-agent sandboxes. Beyond the benchmarks discussed above, additional frameworks strengthen these evaluation dimensions. Red-team corpora like HarmBench [145] and JailbreakBench [29], though primarily LLM-focused, provide adversarial baselines applicable to agent-underlying models. Extending evaluation to realistic web interaction scenarios, dynamic web testbeds such as AgentDojo [61] and WASP [69] assess prompt injection vulnerabilities in web agents. OS-agent sandboxes including ToolEmu [188] and OS-Harm [8] further enable safe evaluation across desktop platforms.
@@ 6.4 Safety Against Adaptive Attacks
6.5 Agentic AI Security in the Physical World
While this survey primarily focuses on software-based agents, physically embodied agents introduce additional and largely unexplored security risks due to their dependence on hardware and direct interaction with the physical environment. (1) Such risks involve adversaries manipulating robot perception through sensor spoofing, including GPS spoofing that redirects vehicles into restricted zones [114, 213], LiDAR laser projection spoofing that injects phantom obstacles triggering emergency braking [21, 200], and adversarial patches on road signs and persons causing threats through misclassification [70, 212]. (2) Ensuring robust and secure operation of physically embodied agents remains a critical challenge, as agent architecture increasingly determines vulnerability. End-to-end foundation models (e.g., RT-2 [271], OpenVLA [115]) map raw sensory inputs directly to motor actions, eliminating defensive separation and prioritizing zero-shot generalization, but without corresponding improvements in adversarial robustness [250]. Addressing this trade-off between generalization and safety, requires new methods to preserve internal defenses and safeguard agents against real-world sensor and hardware attacks. (3) Beyond mobile robotics, LLM-based agents are increasingly deployed in smart home environments for device control and automation [186], yet systematic security evaluation and benchmark in these environments remains underexplored.